Topic:Arithmetic Mean
1. Find the value of b when the arithmetic mean of b+2 and 4b+5 is b+2

Answer by htmentor(1343)

The arithmetic mean (average) is given by ( (b+2) + (4b+5) )/2, and this is equal to b+2:
( (b+2) + (4b+5) )/2 = b+2
Solve for b:
5b + 7 = 2b + 4
b = -1
Check:
(1 + 1)/2 = -1 + 2 = 1